Rocket observations show that dust particles in Earth's upper atmosphere are often electrically charged.

(a) Find the distance separating two dust particles (in m) if each has a charge of +e and the Coulomb force between them has magnitude 1.65 ✕ 10^−14 N.

m

(b) Calculate the mass (in kg) of one of the dust particles if this Coulomb force would accelerate it at 2.60 ✕ 10^8 m/s^2. (In the upper atmosphere, effects from other nearby charges typically result in a small net force and acceleration.)

kg
Answer #1

a)

Coulomb force between the two dust particles is 9192 4TE0/2

Charge on each dust particle is g1 = 92 =-c = +1.6 * 10-19 C

Distance between the dust particles is r

F 1.65 * 10-14 N

91q2

γί99 *109 * (1.6.110 i942 = 1.15+10 r = / 8.99 * 109 * 1.6 10-19)2 1.65 10-14 1.18 10-7, 7n

Distance between the dust particles is 1.18 10m

b)

Acceleration of particle is a 2.60 * 108 m/s2

F=ma

Mass of the dust particle is F 1.65* 10-4 2.60 * 10° m_ = 6.35 * 10-23k